OpenStudy (anonymous):

Arthur is organizing his collection of DVD'S he has 120 altogether The Ratio of comedy films to action to horror is 2:8:5 How many of each type of DVD'S Does he have Comedy.....DVD'S Action......DVD'S Horror.......DVD'S

OpenStudy (anonymous):

2+8+5=15 120/15=8 So he has 8 sets of DVDs, each set consisting of 2 comedies, 8 action films, and 5 horror films.

OpenStudy (anonymous):

so what the answer

OpenStudy (anonymous):

If I have 8 groups of 2 items, how many items do I have?

OpenStudy (anonymous):

Dunno

OpenStudy (anonymous):

It's the same as\[8 \times 2\]